Please solve by addition method:
x - y = 3
-6x + 6y = 17

Answer by rapaljer(4551)

To solve this by the addition method, you need to have the same coefficient for one of the variables, but with opposite signs. In this case, since you have a -y and a +6y, it would be nice to multiply both sides of the first equation by 6, which would give you a -6y and a +6y that will subtract out!

6(x - y) = 6*3
-6x + 6y = 17

6x - 6y = 18
-6x + 6y = 17

Add the two equations together, and you get:
0 + 0 = 35
This is NOT possible, so there is NO SOLUTION. What causes no solution is the fact that the two lines are actually parallel!

R^2 at SCC